OpenStudy (anonymous):

Please help Urgent! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

@matimaticas @Grazes

OpenStudy (anonymous):

Slope intercept form is y=mx+b. Put all the terms without y on one side and keep the term with y on the other side. If there is any constant attached to the y, divide or multiply it to isolate the y. Afterwards, you should know that the slope(in slope intercept form) is m and the y-intercept b.

OpenStudy (anonymous):

Ok but like I have a qestion do I substitute y=mx+b with something in the paragraph?

OpenStudy (anonymous):

@Grazes

OpenStudy (anonymous):

It's not substitution. It's conversion. You have 2x + 3y = 1,200 which is in the Ax+By=C You need to change that into y=mx+b ex. 4x+7y=28 7y= -4x+28 y= -(4/7)x+4

OpenStudy (anonymous):

oh thanks